Your banker has analyzed your company account and has suggested that her bank has a cash management package for you. She suggests that with a concentration banking system, your float can be reduced by three days on average. You, of course, are delighted (you’re not sure why), but you do know your average daily collections amount to $320,000. Your opportunity cost of funds is 9 percent. The bank provides this service for $54,000 plus a compensating balance in your current account of $76,500. (A compensating balance is the amount you are required to maintain interest free at that bank.)

By how much? (Negative answer should be indicated by a minus sign.)

Annual Saving           $
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

Step 1

Finding the cost of funds saved by reducing the float days

=Average daily collection * Float days reduced*cost of funds

=$320,000*3*9%

=$86,400

Step 2

Finding the cost of package

=Bank Service fee+cost of funds on compensating balance

=$54,000 +($76,500*9%)

=$60.885

The package should be accepted as the cost saved by package(step 1) is higher than cost of package.(step 2)

=> Annual savings = =$86,400 - =$60.885

=$25,515
Answer #2

SOLUTION :


Collection in one day = $320000 .

Collection in 3 days = $320000 * 3 = $960000 .

Collection period is reduced by 3 days means cash of $960000 will not be blocked in the system and will be available to the company.


So, 


Saving of interest on this cash-increase in cash availability throughout the year

= 960000 * 0.09

= 86400 ($).


Cost of getting this system 

= Fees of the bank + loss of interest on compensating balance

= 54000 + 76500*0.09

= 60885 ($)


A.


The package offered by the bank is  worth taking since savings are more than the cost. (ANSWER).


B.


The extent of gain

= Savings - Cost 

= 86400 - 60885 

= 25515 ($)  (ANSWER).

answered by: Tulsiram Garg